# Cooperation as Black Box: Conceptual Fluctuation and Diagnostic Tools for Misalignment in MAS

### v1

@@ Abstract
Misalignment in multi-agent systems (MAS) is often treated as a technical failure; yet many such failures originate upstream, during the conceptual design phase, where semantic ambiguity and normative projection take place. This paper identifies a foundational source of interpretive misalignment in MAS: the systemic conflation of cooperation and coordination, and the moral overreading that follows. Using the Rabbit–Duck illusion, we illustrate how perspective-dependent readings of agent behavior can create epistemic instability. To address this, we introduce the Misalignment Mosaic, a diagnostic framework for diagnosing meaning-level misalignment in MAS. It comprises four components: 1. Terminological Inconsistency, 2. Concept-to-Code Decay, 3. Morality as Cooperation, and 4. Interpretive Ambiguity. The Mosaic enables researchers to examine how misalignment arises not only through policy or reward structures but also through language, framing, and design assumptions. While this paper focuses on the specific ambiguity between coordination and cooperation, the Mosaic generalizes to other overloaded concepts in MAS, such as alignment, autonomy, and trust. Rather than define cooperation once and for all, we offer a framework to diagnose meaning itself as a source of misalignment.
@@ 1 Interpretive Ambiguity as Alignment Risk
Consider the following two situations: you notice a stranger struggling to carry a large box and offer help; you then share the physical effort to transport it. Later at home, you and a partner assemble a jigsaw puzzle, coordinating your actions toward a shared goal. What distinguishes cooperation from coordination in these everyday interactions? And how consistently are such distinctions addressed in MAS research?
@@
The inability to clearly separate these concepts reflects a systemic conflation, with real structural consequences. Coordinated behavior is often misread as cooperation, triggering assumptions of moral alignment or shared intent where none exists. This interpretive leap, rooted in the observer rather than the agent, distorts how MAS are evaluated, trusted, and governed, as cooperation surfaces linkages with morality [1]: when coordination is interpreted as cooperation, systems may be evaluated not on what they do, but on what we assume they mean.
@@
Coordinating agents under uncertainty is a core topic in MAS, rooted in distributed systems [2], game theory [3], and machine learning [4]. Similarly, cooperation in multi-agent learning (MAL) intersects disciplines such as economics, social sciences, and evolutionary biology [5]. Within MAS, diverse research traditions, from reinforcement learning (RL) to normative agent design, bring various assumptions about what cooperation entails and how it should be modeled. While terminological flexibility enables methodological innovation, ambiguity between coordination and cooperation introduces misalignment, allowing moral interpretations to influence the design and evaluation of MAS.
@@
As with our opening example, the distinction between coordination and cooperation is often context-dependent, shaped by factors such as the task, environment, and agents’ affordances [6] [7]. Such contextual variability has likely contributed to the interchangeable use of these terms across studies. Yet this ambiguity introduces unintended consequences in MAS research and applications, particularly as other domains build on MAS foundations. Although our focus is on MAS, similar risks arise in emerging Generative AI systems that integrate agent-like structures and decision processes. As these systems scale, ambiguity becomes operational, with human observers projecting intent, values, or goals onto structural outputs. This paper invites MAS researchers to audit how conceptual assumptions migrate into system design (often with normative implications beyond the technical layer).
@@
The coordination/cooperation ambiguity mirrors the Rabbit–Duck illusion (see Figure 1), where perception shifts without changes to the underlying image. Similarly, MAS behavior may be framed as coordination or cooperation depending on the observer’s perspective and assumptions. Then, what initially appears as mere coordination may, upon reinterpretation, reveal underlying cooperative dynamics, just as shifting one’s perception of the image reveals the duck within the rabbit. This ambiguity deepens when viewed through the lens of Braitenberg’s thought experiments [8]: the vehicles operate under simple rules, yet observers frequently project complex intentions onto them [9]. Such perspective-dependent interpretations complicate the design, analysis, and evaluation of MAS, as system behavior becomes linked with the observer’s framing, rather than reflecting only the system’s architecture.
@@
A more intentional effort to distinguish coordination from cooperation is overdue, along with shared terminology, benchmarks [10], and interpretive literacy to scaffold rigorous MAS design and evaluation. To address this, we introduce the Misalignment Mosaic, a framework for diagnosing how semantic instability and interpretive ambiguity migrate from language to system design and ultimately to evaluation failures.
@@
Rather than compress methodological diversity, the Mosaic fosters shared vocabulary to diagnose meaning-related misalignments that precede technical or behavioral failures. Our work complements recent alignment research by focusing on meaning itself, not only how agents behave, but how their behavior is framed, trusted, and morally interpreted in MAS. This semantic gap remains structurally unaddressed in MAS, despite its growing relevance in public-facing systems.
@@
The term mosaic reflects that these misalignments occur across shapes, scales, and contexts, collectively forming a complex interpretive heterogeneity — a phenomenon of alignment magnification [11]. In its current structure, the Mosaic comprises four diagnostic components, detailed in Section 3 and shown in Figure 1. Our main contributions are:
  1. 1.
@@
    We identify three interconnected drivers of interpretive misalignment in MAS: (1) coordination–cooperation ambiguity, where these terms are used interchangeably, blurring conceptual boundaries; (2) conceptual fluctuation, where key terms like cooperation acquire shifting meanings across system architecture, agent behavior, evaluation, and observer framing; and (3) semantic instability, where the anchoring of these terms becomes fragile or inconsistent, allowing meaning to drift across subfields, implementations, or observer perspectives. These dynamics systematically enable normative overreadings, where coordination can be misread as cooperation, and cooperation as moral alignment, producing interpretive, design-level, and evaluative risks.
  2. 2.
@@
    We introduce the Misalignment Mosaic: a framework for diagnosing meaning-level misalignment in MAS. The Mosaic consists of four diagnostic components: (1) Terminological Inconsistency, (2) Concept-to-Code Decay, (3) Morality as Cooperation, and (4) Interpretive Ambiguity. Together, these components help trace how misalignment arises not only from agent architectures or policies but from linguistic, epistemic, and framing assumptions embedded in MAS.
@@
Variations in how coordination and cooperation are defined influence theoretical models, experimental design, and interdisciplinary communication. These inconsistencies are particularly problematic in multidisciplinary and multilingual research settings, where different conceptual models silently shape interpretation and evaluation. Rather than prescribing fixed definitions, we argue for diagnosing semantic instability and interpretive ambiguity as structural risks, which intensify as MAS scale into institutional and societal domains, where ambiguity itself becomes operational and must be surfaced before alignment can be assessed or governed.
Figure 1: Left: The Rabbit–Duck illusion, illustrating how observer perspective shapes classification in MAS (is behavior seen as coordination or cooperation)? Misalignment Mosaic’s motivation at the Center, and at Right, its visual structure (a diagnostic framework for tracing semantic and interpretive misalignment).
@@ 2.1 Semantic Instability and Everyday Settings
Familiar settings show the similarity between coordination and cooperation: both structure how humans divide labor, respond to requests, and signal intent, e.g., “Thanks for your cooperation,” or “Let’s coordinate on this.” Yet when imported into MAS design or evaluation frameworks, their meanings become unstable: is cooperation defined by agent behavior, internal intent, task structure, or outcome? What distinguishes cooperation from strategic alignment, sacrifice, or simple compliance? Standard definitions provide little clarity: coordination is often framed as “the process of organizing people or groups so that they work together properly and well” or “the harmonious functioning of parts for effective results” [12], while cooperation implies shared effort or helpfulness [13].
@@
Why does it matter? Semantic instability can propagate misunderstandings and micro-interpretations as systems scale in complexity, especially when more advanced frameworks are built on top of them. If key terms are loosely defined at lower levels, the resulting inconsistencies will spread to more sophisticated applications. In fact, Barad [14] argues that matter and meaning are inextricably linked, emphasizing that scientific inquiry must account for the inseparability of epistemological, ontological, and ethical considerations.
@@
In fact, Langer et al. [15] investigate if terminological differences affect people’s 1) Perceptions of properties of algorithmic decision-making systems, 2) Evaluations of systems in application contexts, and 3) The robustness and replicability of human-computer-interaction research. Their results emphasize the importance of carefully selecting terminology, as it can significantly bias users’ interpretations and judgments of algorithmic systems – the Citizen Science[16] examines similar ideas. Similarly, aiming to build systems that can describe what they perceive using the users’ terms, [17] claims that object recognition systems should follow a process that resembles the human when thinking of an object associated with a certain concept.
@@
In MAS, agents may coordinate without explicit cooperation or cooperate without centralized coordination. Cooperation itself is overloaded: spanning strategic alliances, sacrificial gestures, or even tacit collusion. Related concepts, such as collaboration, assistance, opposite coordination, required cooperation [18], collusion, coalition, cooperative-competitive dynamics [19], and advantage alignment [20], further complicate the landscape. Each variation implies different agent behaviors, moral interpretations, and normative assessments, underscoring the importance of clarity in MAS design and evaluation. Similar issues extend beyond MAS; in fact, Mesly et al.  [21] analyze how economic academics inconsistently define “predation” and sometimes even erroneously or too narrowly.
@@
To surface issues at play, we raise the following questions for reflection:
@@
    Is coordination a structural prerequisite for cooperation[22], or are the two orthogonal? (Do different definitions make that answer fluctuate?)
@@
    Does explicit instruction or internally inferred “goal” to cooperate with others assume coordination?
@@
    Is collaboration a subset of cooperation? If so, what does that imply about hierarchical structures across different ways to define “cooperation”?
@@
    Can cooperation be evaluated without invoking moral projection (more in section 3)?
@@
Malone and Crowston [23] define coordination as: “the act of managing interdependencies between activities performed to achieve a goal”. Building on that, agent coordination can be understood as the “ability to manage the interdependencies of activities between agents” [24]. In contrast, agent cooperation involves the “process used for an agent to voluntarily enter a relationship with another to achieve a system derived goal” [24]. Coordination often highlights agents’ inter-dependencies management or task-level alignment, whereas cooperation often invokes common goals, altruistic behavior, and even moral norms (the “Morality-as-Cooperation” Theory). This distinction is further illustrated by Jensen et al. [25], who examine how people use accents of unfamiliar individuals to make social decisions through economic games that measure interpersonal trust, generosity, and coordination. Their results suggest that humans use ethnic markers of unfamiliar individuals to coordinate behavior rather than to cooperate.
Upon closer examination of the MAS literature, it becomes clear that the definition of cooperation varies significantly. For example, Tan [26] contrasts independent learners with cooperative agents, defining cooperation as agents sharing episodes, learned policies, or instantaneous information, whereas others may take a different approach. As is the case with [27] (following [28]), “cooperative robots” and “fully cooperative MAS learning algorithms” are settings where agents share common interests or the same utility function. In such cases, there is a correspondence between each agent’s achievement and the group’s, and therefore, the learning goal is defined as maximizing the common discounted return. Or, as [28] puts it, fully cooperative MAS in which we assume that it is possible to set a common coordination mechanism and that agents do not have a reason to deliberate strategically. These variations highlight the terminological inconsistencies in MAS research, where cooperation can mean anything from simple information sharing to fully shared objectives and policies. Addressing these distinctions is essential for developing a coherent theoretical foundation for MAS.
While the empirical studies above tackle different challenges in MAS, we argue that three interconnected drivers (coordination/cooperation ambiguity, conceptual fluctuation, and semantic instability) must be addressed before empirical methodologies can be effectively compared. Those issues lead to terminological inconsistencies in MAS research, resulting in conceptual fragmentation and making it difficult to establish common frameworks and methodologies.
We identify a few factors that compound the three interconnected drivers by making them harder to detect, evaluate, or resolve:
@@
    Lack of uniformity in defining environments. MAS environments are often abstracted or specified inconsistently [29], which obscures whether observed cooperation arises from agent design, environmental affordances, or emergent properties.
@@
    Conflicting research agendas. Hinder the evaluation and comparability of results [30], making it difficult to compare claims about cooperative behavior.
@@
    Architectural-task Inconsistency. Some approaches define cooperation in terms of agent intent or architecture, while others define it purely in terms of task outcomes. For instance, in the Prisoner’s Dilemma, even rule-based agents are said to “cooperate” if their behavior aligns with the game’s cooperative payoff [31]. This further blurs the boundary between coordination and cooperation in agent interpretation.
The SMAC environment [32] provides a set of benchmark tasks for cooperative MARL, offering standardized challenge scenarios and best-practice guidelines. PettingZoo [33] extends these efforts with a multi-agent environment library and a universal API to accelerate MARL research. More recently, Cross-environment Cooperation [34] explores zero-shot multi-agent coordination. Earlier work by Fulda and Ventura [35] highlighted the importance of predicting and preventing coordination problems in cooperative Q-learning systems. Our work complements such initiatives by addressing a preceding and often overlooked layer: the need for consistency at the meaning level. Instability in terms like cooperation, coordination, and alignment across technical benchmarks risks reinforcing interpretive ambiguity rather than mitigating it.
Our goal is not to evaluate or craft definitions of coordination and cooperation but to highlight the urgency of the issues. Addressing this involves a shared conceptual vocabulary, a task that requires a concerted multidisciplinary effort because cooperation, as it manifests in MAS, crosses conceptual boundaries. It is not only a matter of policy or control, but of intention modeling, value projection, and interpretive inference, all of which are drawn from fields such as philosophy, cognitive science, and ethics. (This work acknowledges its own positioning within these evolving conceptual boundaries.)

### v2

# Cooperation as a Black Box: Conceptual Fluctuation and Diagnostic Tools for Misalignment in MAS

@@ Abstract
Misalignment in Multi-Agent Systems (MAS) is frequently treated as a technical failure. Yet, issues may arise from the conceptual design phase, where semantic ambiguity and normative projection occur. The Rabbit–Duck illusion illustrates how perspective-dependent readings of agent behavior, such as the conflation of cooperation/coordination, can create epistemic instability; e.g., coordinated agents in cooperative Multi-Agent Reinforcement Learning (MARL) benchmarks being interpreted as morally aligned, despite being optimized for shared utility maximization only. Motivated by three drivers of meaning-level misalignment in MAS (coordination–cooperation ambiguity, conceptual fluctuation, and semantic instability), we introduce the Misalignment Mosaic: a framework for diagnosing how misalignment emerges through language, framing, and design assumptions. The Mosaic comprises four components: 1. Terminological Inconsistency, 2. Interpretive Ambiguity, 3. Concept-to-Code Decay, and 4. Morality as Cooperation. Building on insights from the Morality-as-Cooperation Theory, we call for consistent meaning-level grounding in MAS to ensure systems function as intended: technically and ethically. This need is particularly urgent as MAS principles influence broader Artificial Intelligence (AI) workflows, amplifying risks in trust, interpretability, and governance. While this work focuses on the coordination/cooperation ambiguity, the Mosaic generalizes to other overloaded terms, such as alignment, autonomy, and trust.
@@ 1 Interpretive Ambiguity as Alignment Risk
Consider the following two situations: you notice a stranger struggling to carry a large box and offer help; you then share the physical effort to transport it to a goal location. Later at home, you and a partner assemble a jigsaw puzzle, coordinating your actions toward a shared goal. What distinguishes cooperation from coordination in these everyday interactions? And how consistently are such distinctions, or ambiguity, addressed in MAS research?
@@
Challenges in consistently separating these terms [^1] We anchor “concept” to a word’s associated abstract ideas [concept_def], whereas “term” to its more precise or technical meaning [term_def]. reflects a systemic conflation, with real structural consequences. Coordinated behavior is often misread as cooperation, leading to assumptions of moral alignment or shared intent where none exist. This interpretive jump, rooted in the observer rather than the agent, distorts how MAS are evaluated, trusted, and governed, as cooperation has ties to morality [tomasello2013origins]: when coordination is interpreted as cooperation, systems may be evaluated on what we assume they mean instead of on what they do.
@@
Coordinating agents under uncertainty is a core topic in MAS, rooted in distributed systems [lynch1996distributed], game theory [chakravarty2014course], and machine learning [lanctot2017unified]. Similarly, cooperation in Multi-Agent Learning intersects disciplines such as economics, social sciences, and evolutionary biology [du2023review]. Within MAS, diverse research traditions (e.g., Reinforcement Learning (RL), normative agent design) bring various assumptions about what cooperation entails and how it should be modeled. While terminological flexibility enables methodological innovation, the coordination/cooperation ambiguity is a tricky case because it introduces moral interpretations into the design and evaluation of MAS.
@@
As with our opening example, the distinction between coordination/cooperation is often context-dependent, shaped by factors such as the task, environment, and agents’ affordances [gibson1966senses] [gibson2014ecological]. Contextual variability has likely contributed to the interchangeable use of these terms across studies; however, this ambiguity introduces unintended consequences in MAS research and applications, particularly as other domains build upon MAS foundations. In fact, similar risks arise in emerging Generative AI systems that integrate agent-like structures and decision processes. As these systems scale, ambiguity becomes operational, with human observers projecting intent, values, or goals onto structural outputs. Thus, this paper invites MAS researchers to audit how semantic ambiguity and conceptual assumptions migrate into system design and deployment (often with normative implications beyond the technical layer).
@@
The coordination/cooperation ambiguity mirrors the Rabbit– Duck illusion (see Figure 1), where perception shifts without changes to the underlying image. Similarly, MAS behavior may be framed as coordination or cooperation depending on the observer’s perspective and assumptions. Then, what initially appears as “coordination” may, upon reinterpretation, reveal underlying cooperative dynamics, just as shifting one’s perception of the image reveals the duck within the rabbit. This ambiguity deepens when viewed through the lens of Braitenberg’s thought experiments [braitenberg1986vehicles]: the vehicles operate under simple rules, yet observers easily project complex intentions onto them (see [prescott2019synthetic], [shaikh2020braitenberg], and  [Vehicles2023] for more perspectives on synthetic psychology and the vehicles). Such perspective-dependent interpretations complicate the design, analysis, and evaluation of MAS, as system behavior becomes linked with the observer’s framing, rather than reflecting only the system’s architecture.
@@
Therefore, we claim that a more intentional effort to distinguish coordination from cooperation is overdue, along with shared terminology, benchmarks [hanks1993benchmarks], and interpretive literacy to scaffold rigorous MAS design and evaluation. To show how the problem extends beyond technical misalignment, we introduce the Misalignment Mosaic: a framework for diagnosing how semantic and interpretive ambiguity migrate from language to system design and ultimately to evaluation failures.
@@
Rather than compressing methodological diversity, the Mosaic fosters shared vocabulary and practices. It complements recent alignment research by focusing on meaning itself, not only on how agents behave, but also on how their behavior is framed and morally interpreted in MAS. This semantic gap remains structurally unaddressed in MAS, despite its growing relevance in public-facing systems.
@@
The term “mosaic” illustrates that MAS misalignments occur across shapes, scales, and contexts, collectively forming a complex interpretive heterogeneity — a phenomenon of alignment magnification [dung2023current]. In its current structure, the Misalignment Mosaic comprises four diagnostic components, detailed in Section 3 and shown in Figure 1. Our main contributions are:
@@
(1) We identify three interconnected drivers of interpretive misalignment in MAS: a) Coordination/cooperation ambiguity, where these terms are used interchangeably, blurring conceptual boundaries; b) Conceptual fluctuation, where key terms like cooperation acquire shifting meanings across system architecture, agent behavior, evaluation, and observer framing; and c) Semantic instability, where the anchoring of key terms becomes fragile or inconsistent, allowing meaning to drift across subfields, implementations, or observer perspectives. These dynamics systematically enable normative overreadings, in which coordination can be misread as cooperation and cooperation as moral alignment, producing interpretive, design-level, evaluative, and ultimately ethical risks.
@@
(2) We introduce the Misalignment Mosaic: a framework for diagnosing meaning-level misalignment in MAS. It comprises four components: 1. Terminological Inconsistency, 2. Interpretive Ambiguity, 3. Concept-to-Code Decay, and 4. Morality as Cooperation. Together, these components help trace how misalignment arises not only from agent architecture or policy but from linguistic, epistemic, and framing assumptions embedded in MAS.
Variations in how coordination and cooperation are defined influence theoretical models, experimental design, and interdisciplinary communication. Inconsistencies are particularly problematic in multidisciplinary and multilingual research settings, where different conceptual models silently shape interpretation and evaluation. Rather than prescribing fixed definitions, we argue that diagnosing semantic instability and interpretive ambiguity as structural risks is appropriate, as these risks intensify as MAS scale into institutional and societal domains, where ambiguity itself becomes operational and must be recognized before alignment can be assessed or governed.
@@
Figure 1: Left: The Rabbit–Duck illusion (made famous by Wittgenstein [wittgenstein]), illustrating how observer perspective influences MAS (e.g., is behavior seen as coordination or cooperation?). Misalignment Mosaic’s three motivations at the Center, and at Right, a visual depiction (a diagnostic framework for tracing semantic and interpretive misalignment).
@@ 2.1 Semantic Instability and Everyday Settings
Familiar settings exemplify the similarity between coordination and cooperation. Both structure how humans divide labor, respond to requests, and signal intent: “Thanks for your cooperation,” or “Let’s coordinate on this.” Yet when imported into MAS design or evaluation frameworks, their meanings become unstable: is cooperation defined by agent behavior, internal intent, task structure, or outcome? What distinguishes cooperation from strategic alignment, sacrifice, or simple compliance? Standard definitions provide little clarity: coordination is often framed as “the process of organizing people or groups so that they work together properly and well” or “the harmonious functioning of parts for effective results” [coordination_def], while cooperation implies shared effort or helpfulness [coop_def].
@@
Why does it matter? Semantic instability can propagate misunderstandings and normative projection as systems scale in complexity, especially when more advanced frameworks are built on top of them. If key terms are variously defined at lower levels, the resulting inconsistencies will spread to more sophisticated applications. In fact, Barad [barad2007meeting] argues that matter and meaning are inextricably linked, emphasizing that scientific inquiry must account for the inseparability of epistemological, ontological, and ethical considerations.
@@
Indeed, Langer et al. [langer2022look] investigate if terminological differences affect people’s 1) Perceptions of properties of algorithmic decision-making systems, 2) Evaluations of systems in application contexts, and 3) The robustness and replicability of human-computer-interaction research. Their results emphasize the importance of carefully selecting terminology, as it can significantly bias users’ interpretations and judgments of algorithmic systems – the Citizen Science[eitzel2017citizen] examines similar ideas. Similarly, aiming to build systems that can describe what they perceive using the users’ terms, Erculiani and colleagues [erculiani2023egocentric] claim that object recognition systems should follow a process that resembles the human when thinking of an object associated with a certain concept.
@@
In MAS, agents may coordinate without explicit cooperation or cooperate without centralized coordination. Cooperation itself is overloaded: spanning strategic alliances, sacrificial gestures, or even tacit collusion. Related terms, such as collaboration, assistance, opposite coordination, required cooperation [zhang2016formal], collusion, coalition, cooperative-competitive dynamics [tuyls2012multiagent], and advantage alignment [duque2024advantage], further complicate the landscape. Each variation implies different agent behaviors, moral interpretations, and normative assessments, underscoring the importance of clarity in MAS design and evaluation. Similar issues extend beyond MAS; in fact, Mesly et al.  [mesly2022terminology] analyze how economic academics inconsistently define “predation” and sometimes even erroneously or too narrowly.
@@
To surface other issues at play, we raise the following questions for reflection:
@@
    Is coordination a structural prerequisite for cooperation[huhns1999multiagent]? (Do different definitions change the answer?)
@@
    Do explicit instruction and/or internally inferred goal to cooperate with others presuppose coordination?
@@
    Can cooperation be evaluated without invoking normative readings and moral projection (more in Section 3)?
@@
    Is collaboration a subset of cooperation? If so, what does that imply about hierarchical structures across different definitions of cooperation?
We do not answer these questions; on the contrary, we hope it helps illustrate the relevance of meaning-level misalignment and how it can spread to technical consequences.
@@
Malone and Crowston [malone1990coordination] define coordination as “the act of managing interdependencies between activities performed to achieve a goal”. Building on that, agent coordination can be understood as the “ability to manage the interdependencies of activities between agents” [consoli2007link]. In contrast, agent cooperation involves the “process used for an agent to voluntarily enter a relationship with another to achieve a system derived goal” [consoli2007link]. Coordination often highlights agents’ inter-dependencies management or task-level alignment, whereas cooperation often invokes common goals, altruistic behavior, and even moral norms (the “Morality-as-Cooperation” Theory). This distinction is further illustrated by Jensen et al. [jensen2015testing], who examine how people use accents of unfamiliar individuals to make social decisions through economic games that measure interpersonal trust, generosity, and coordination. Their results suggest that humans use ethnic markers of unfamiliar individuals to coordinate behavior rather than to cooperate.
Upon closer examination of the MAS literature, it becomes clear that the definition of cooperation varies significantly. For example, Tan [tan1993multi] contrasts independent learners with cooperative agents, defining cooperation as agents sharing episodes, learned policies, or instantaneous information, whereas others may take a different approach. As in [matignon2012independent] (following [boutilier1999sequential]), “cooperative robots” and “fully cooperative MAS learning algorithms” are settings where agents share common interests or the same utility function. In such cases, there is a correspondence between each agent’s achievement and the group’s, and therefore, the learning goal is defined as maximizing the common discounted return. Or, as [boutilier1999sequential] puts it, fully cooperative MAS in which we assume that it is possible to set a common coordination mechanism and that agents do not have a reason to deliberate strategically. These variations highlight the terminological inconsistencies in MAS research, where cooperation can mean anything from simple information sharing to fully shared objectives and policies. Addressing these distinctions is essential for developing a coherent theoretical foundation for MAS.
While the empirical studies above tackle different challenges in MAS, we argue that three interconnected drivers (coordination/cooperation ambiguity, conceptual fluctuation, and semantic instability) must be addressed before empirical methodologies can be effectively compared. Those issues lead to terminological inconsistencies in MAS research, resulting in conceptual fragmentation and making it difficult to establish common frameworks and methodologies. We identify a few factors that compound the three interconnected drivers by making them harder to detect, evaluate, or resolve:
@@
    Lack of uniformity in defining environments. MAS environments are often abstracted or specified inconsistently [weyns2006environments], which obscures whether observed cooperation arises from agent design, environmental affordances, or emergent properties.
@@
    Conflicting research agendas. Hinder the evaluation and comparability of results [shoham2007if], making it difficult to compare claims about cooperative behavior.
@@
    Architectural-task inconsistency. Some approaches define cooperation in terms of agent intent or architecture, while others define it purely in terms of task outcomes. For instance, in the Prisoner’s Dilemma, even rule-based agents are said to “cooperate” if their behavior aligns with the game’s cooperative payoff [axelrod1981evolution]. This further blurs the boundary between coordination and cooperation in agent interpretation.
The SMAC environment [samvelyan2019starcraft] provides a set of benchmark tasks for cooperative MARL, offering standardized challenge scenarios and best-practice guidelines. PettingZoo [terry2021pettingzoo] extends these efforts with a multi-agent environment library and a universal API to accelerate MARL research. More recently, Cross-environment Cooperation [jha2025cross] explores zero-shot multi-agent coordination. Earlier work by Fulda and Ventura [fulda2007predicting] highlighted the importance of predicting and preventing coordination problems in cooperative Q-learning systems. Our work complements such initiatives by addressing a preceding and often overlooked layer: the need for consistency at the meaning level. Instability in terms like cooperation and coordination across technical benchmarks risks reinforcing interpretive ambiguity rather than mitigating it.
Sun et al. [sun2025multi] survey coordination research and propose a unified three-component framework for coordination in sequential decision-making: system-level performance evaluation, social choice on who to coordinate with, and how to coordinate. While these models advance technical coordination, our work foregrounds an overlooked dimension — how semantic ambiguity, conceptual fluctuation, and moral overreadings surrounding cooperation complicate the design and evaluation of MAS.
Our goal is not to evaluate or craft definitions of coordination and cooperation but to highlight the urgency of the issues. Addressing this requires a shared conceptual vocabulary, a task that calls for a concerted multidisciplinary effort because cooperation, as it manifests in MAS, crosses conceptual boundaries. It is not only a matter of policy or control, but of intention modeling, value projection, and interpretive inference, all of which are drawn from fields such as philosophy, cognitive science, and ethics. (This work acknowledges its own positioning within these evolving conceptual boundaries.)